Find a function y of x such that


9yy'=x and y(9)=10

1
Expert's answer
2022-03-13T17:34:51-0400

The equation "9yy'=x" is equivalent to "9ydy=xdx." It follows that "\int 9ydy=\int xdx," and hence the general solution is "\frac{9}2y^2=\frac{1}2x^2+C" which is equivalent to "9y^2=x^2+C."


Since "y(9)=10," we get that "9\cdot 10^2=9^2+C," and hence "C=819."

We conclude that the particular solution is "9y^2=x^2+819."